OpenStudy (anonymous):

Given the equations y = x2 - 4x - 5 and y + x = -1, one point that satisfies both equations is: [1] (-2,1) [2] (4,-5) [3] (2,-9) [4] (5,0)

OpenStudy (gorv):

y = x2 - 4x - 5 y + x = -1.........implies y=-x-1 -x-1=x2-4x-5 0=x2-4x-5+x+1 0=x2-3x-4 0=x2-4x+x-4 0=x(x-4)+1(x-4) 0=(x+1)(x-4) x=-1 & x=4

OpenStudy (gorv):

now put x in y+x=-1 for x=-1 y-1=-1 y=0 for x=4 y+4=-1 y=-5

OpenStudy (gorv):

(4,-5)

OpenStudy (gorv):

@chitij

OpenStudy (anonymous):

than gorv

OpenStudy (gorv):

ans is (4,-5)

OpenStudy (gorv):

@chitij